[Hallc_running] [New Logentry] RC Update Friday Feb 28

brads at jlab.org brads at jlab.org
Fri Feb 28 13:30:03 EST 2020

Logentry Text:
<p><strong><a class="new" title="RC Meeting and Run Plan, Friday, February 28th 2020 (page does not exist)" href="https://hallcweb.jlab.org/wiki/index.php?title=RC_Meeting_and_Run_Plan,_Friday,_February_28th_2020&action=edit&redlink=1">RC Meeting and Run Plan, Friday, February 28th 2020</a></strong></p><p><strong>RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room </strong> <a class="external text" href="https://bluejeans.com/861439920" rel="nofollow">Bluejeans Connection</a></p><p><strong> Friday 28th Feb. 2020. </strong></p><h1><span id="General_updates" class="mw-headline">General updates</span></h1><ul><li>Problems with 1L07, 1L17 throughout day shift (trippy beam, made Moller difficult, etc)</li><li>Lost beam mid-Swing and had extended down through OWL (issues in the NL, same cavities)<ul><li><a class="external free" href="https://cebaf.jlab.org/dtm/open-events?event_id=6195" rel="nofollow">https://cebaf.jlab.org/dtm/open-events?event_id=6195!
 </a></li><li>Found that 1L20 had been misconfigured a day or two ago and was 145deg out of phase, now corrected</li><li>Issues with 1L07-7 and 1L07-8 looked pretty grim in late swing, but seem to have been recoverable(?)</li></ul></li><li>Hall C target ion chamber tripped with no beam in the machine <a class="external autonumber" href="https://logbooks.jlab.org/entry/3793622" rel="nofollow">[1]</a></li><li>Moller results from Thursday (85.0%) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3793398" rel="nofollow">[2]</a></li><li>MCC took a few hours this morning to install an 'Easy Button' software patch and to chase some of the issues from overnight<ul><li>We took an opportunistic access to take a look at a few items:<ul><li>Swap USB-ethernet bridge with our spare -- Done (an attempt to address the coil current readback issues w/ LabView)</li><li>Looked at NMR and EPR hardware connections/settings (<a class="external free" href="https://logbooks.jlab.o!
 rg/entry/3793821" rel="nofollow">https://logbooks.jlab.org/ent!
 ry/3793821</a>)<ul><li>Found and corrected an internal vs. external trigger source setting switch with the EPR setup</li><li>This will also help (maybe solve?) the issue with the transverse NMR readback issues <a class="external autonumber" href="https://logbooks.jlab.org/entry/3793881" rel="nofollow">[3]</a></li><li>No 'smoking gun' found that could explain the large polarization drop on 2/24 though</li></ul></li></ul></li></ul></li></ul><h2><span id="Notes_for_the_Future" class="mw-headline">Notes for the Future</span></h2><ul><li>Beam studies to understand what is going on with the machine are planned for Tuesday DAY shift<ul><li>Details being worked out, but it'll be a full 7--8 hour down (for us)</li><li>EPR and pNMR measurements/work are planned during this downtime</li><li>There are portions of the Beam Studies program that would involve beam delivery into Hall C (no target), so limited opportunity for Access only</li></ul></li><li>Injector spot move has been pushed !
 off for at least another week (2--3 hour down)<ul><li>Not much beam over the last several days...</li><li>Hope is to push it off to the Hall C down that starts March 13, and then just do a reactivation that would carry us through May 6<ul><li>Not sure how it will work out, but goal will be to minimized downs and maximize integrated current to Hall(s)</li></ul></li></ul></li></ul><ul><li>1 hour run with target out for measure of the background just from the beryllium windows (<em>Filler</em> for next spin-up)</li></ul><ul><li><strong>Target coil rotation planned for March 13--20</strong><ul><li>Nominal 1-week downtime in Hall C to rotate target coils, install new cell, and transition to d2n program</li><li>Status Update for target cell <em>Austin</em>:<ul><li>Wall and window thicknesses done</li><li>Only 'Reduced set' of density measurements available now, will continue measurements this week (issues with laser at UVa?)</li><li>It would be best for it to arrive at JLab on Ma!
 rch 10 (at the latest) to allow locals to mount cell in oven plate befo!
 re they get involved with target work in Hall.</li></ul></li><li>Lots of preplanning to be completed (RadCon, Techs, S&A, Target group, etc)</li><li>RW-II training for JP, Arun, Brad (Complete the training documentation and take the RW-II test if you haven't already done so)</li></ul></li></ul><h2><span id="Pending_problems" class="mw-headline">Pending problems</span></h2><ul><li>Beam energy shifts between ARC1:p and INJ:p impacted beam polarization to B/C? (Started 2/18 through 2/28) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3793295" rel="nofollow">[4]</a><ul><li><strong>Resolved</strong> <a class="external autonumber" href="https://logbooks.jlab.org/entry/3793800" rel="nofollow">[5]</a></li></ul></li><li>3C20 BPM noise issue <a class="external free" href="https://logbooks.jlab.org/entry/3793274" rel="nofollow">https://logbooks.jlab.org/entry/3793274</a><ul><li><strong>Resolved</strong>: BPMs found to be in FFB mode, now back in normal mode (!
 Jay)</li></ul></li><li>Persistent BLM & ion chamber trips<ul><li>General activation, weird BLM issues (<a class="external free" href="https://logbooks.jlab.org/entry/3792326" rel="nofollow">https://logbooks.jlab.org/entry/3792326</a>), Ad-hoc ion chamber threshold adjustments (<a class="external free" href="https://logbooks.jlab.org/entry/3792095" rel="nofollow">https://logbooks.jlab.org/entry/3792095</a>)</li><li><strong>NOTE</strong>: Shift crew should <em>NOT</em> give authorization to MCC change ion chamber thresholds. If they ask, tell Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)</li></ul></li></ul><h2><span id="Fixed_problems" class="mw-headline"><br></span></h2>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3793897
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200228/a31a3aa6/attachment.html>

More information about the Hallc_running mailing list